# Time Series Foundation Models: Benchmarking Challenges and Requirements

### v1

@@ Appendix C TSFM Data Lineage
Dataset Source Domain Frequency Chronos TimesFM Moirai Moirai-MoE Time-GPT ForecastPFN MOMENT Lag-Llama GPT4TS Timer ZSisionTS TTM LLMTime Time-LMM UniTime Time-MoE
LargeST LargeST (Liu et al., 2023a) Transport 5T P P P
Beijing Subway LibCity (Wang et al., 2023a) Transport 30T P P P
HZMetro LibCity (Wang et al., 2023a) Transport 15T P P P
Loop Seattle LibCity (Wang et al., 2023a) Transport 5T P P P
Loop Seattle LibCity (Wang et al., 2023a) Transport H
Loop Seattle LibCity (Wang et al., 2023a) Transport D
Los-Loop LibCity (Wang et al., 2023a) Transport 5T P P P P
M_DENSE LibCity (Wang et al., 2023a) Transport H
M_DENSE LibCity (Wang et al., 2023a) Transport D ZS
PEMS Bay LibCity (Wang et al., 2023a) Transport 5T P P P P
PEMS03 LibCity (Wang et al., 2023a) Transport 5T P P ZS P P
PEMS04 LibCity (Wang et al., 2023a) Transport 5T P P ZS P P
PEMS07 LibCity (Wang et al., 2023a) Transport 5T P P ZS P P
PEMS08 LibCity (Wang et al., 2023a) Transport 5T P P T/T ZS P P
Q-Traffic LibCity (Wang et al., 2023a) Transport 15T P P P P P
SHMetro LibCity (Wang et al., 2023a) Transport 15T P P P P
SZ-Taxi LibCity (Wang et al., 2023a) Transport 15T P P P
SZ-Taxi LibCity (Wang et al., 2023a) Transport H P
Rotterdam LibCity (Wang et al., 2023a) Transport 2T P P
Beijing Air Quality LOTSA_Others (Woo et al., 2024) Nature H P P P P P
CDC Fluview ILINet LOTSA_Others (Woo et al., 2024) Healthcare W P P ZS T/T T/T ZS T/T T/T P
CDC Fluview WHO NREVSS LOTSA_Others (Woo et al., 2024) Healthcare W P P ZS T/T T/T P
China Air Quality LOTSA_Others (Woo et al., 2024) Nature H P P P P
Favorita Sales LOTSA_Others (Woo et al., 2024) Sales D P P P P
Favorita Transactions LOTSA_Others (Woo et al., 2024) Sales D P P P
GoDaddy LOTSA_Others (Woo et al., 2024) Econ/Fin M P P
KDD Cup 2022 LOTSA_Others (Woo et al., 2024) Energy 10T P P P
Project Tycho LOTSA_Others (Woo et al., 2024) Healthcare W P P P
Residential Load Power LOTSA_Others (Woo et al., 2024) Energy T P P P
Residential PV Power LOTSA_Others (Woo et al., 2024) Energy T P P P
@@
Solar LSTNet (Lai et al., 2017) Energy 10T
Solar LSTNet (Lai et al., 2017) Energy 5T P
Solar LSTNet (Lai et al., 2017) Energy H P ZS ZS ZS
Solar LSTNet (Lai et al., 2017) Energy D
Solar LSTNet (Lai et al., 2017) Energy W-FRI
Exchange Rate LSTNet (Lai et al., 2017) Econ/Fin D ZS ZS T/T ZS ZS T/T P
Hierarchical Sales (M5 Competition) Mancuso et al. (2020) Sales D P
Hierarchical Sales (M5 Competition) Mancuso et al. (2020) Sales W-WED
Australian Electricity Demand Monash (Godahewa et al., 2021) Energy 30T ZS ZS T/T T/T P P P ZS P ZS P
Bitcoin Monash (Godahewa et al., 2021) Econ/Fin D ZS T/T T/T P P ZS P ZS P
Car Parts Monash (Godahewa et al., 2021) Sales M T/T T/T P ZS P
CIF 2016 Monash (Godahewa et al., 2021) Econ/Fin M ZS ZS T/T T/T P ZS ZS P
COVID Deaths Monash (Godahewa et al., 2021) Healthcare D ZS ZS T/T T/T P ZS ZS P
Covid Mobility Monash (Godahewa et al., 2021) Transport D T/T T/T P
Elecdemand Monash (Godahewa et al., 2021) Energy 30T T/T T/T P
Extended Web Traffic Monash (Godahewa et al., 2021) Web D T/T T/T P P P
FRED MD Monash (Godahewa et al., 2021) Econ/Fin M ZS ZS T/T T/T P ZS ZS P
Hospital Monash (Godahewa et al., 2021) Healthcare M ZS ZS T/T T/T P ZS ZS P
Kaggle Web Traffic Weekly Monash (Godahewa et al., 2021) Web W P P P P
KDD Cup 2018 Monash (Godahewa et al., 2021) Nature H T/T T/T T/T P P P ZS P P
KDD Cup 2018 Monash (Godahewa et al., 2021) Nature D
London Smart Meters Monash (Godahewa et al., 2021) Energy 30T T/T P P P P P P P
M1 Monthly Monash (Godahewa et al., 2021) Econ/Fin M ZS T/T T/T P ZS P
M1 Quarterly Monash (Godahewa et al., 2021) Econ/Fin Q ZS P P P P
M1 Yearly Monash (Godahewa et al., 2021) Econ/Fin Y ZS P P P P
M3 Monthly Monash (Godahewa et al., 2021) Econ/Fin M ZS T/T T/T ZS T/T ZS P
M3 Other Monash (Godahewa et al., 2021) Econ/Fin Q T/T T/T T/T ZS
M3 Quarterly Monash (Godahewa et al., 2021) Econ/Fin Q ZS P P ZS T/T P
M3 Yearly Monash (Godahewa et al., 2021) Econ/Fin Y ZS P P ZS T/T P
M4 Daily Monash (Godahewa et al., 2021) Econ/Fin D T/T P T/T T/T P T/T ZS P
M4 Hourly Monash (Godahewa et al., 2021) Econ/Fin H T/T P T/T T/T P T/T ZS P
M4 Monthly Monash (Godahewa et al., 2021) Econ/Fin M T/T P T/T T/T ZS T/T ZS P
M4 Quarterly Monash (Godahewa et al., 2021) Econ/Fin Q-DEC ZS P P P ZS T/T P
M4 Weekly Monash (Godahewa et al., 2021) Econ/Fin W-SUN T/T P T/T T/T T/T ZS P
M4 Yearly Monash (Godahewa et al., 2021) Econ/Fin A-DEC ZS P P P ZS T/T P
NN5 Daily Monash (Godahewa et al., 2021) Econ/Fin D ZS ZS T/T T/T P ZS P ZS P
NN5 Weekly Monash (Godahewa et al., 2021) Econ/Fin W ZS ZS T/T T/T P ZS ZS
Oikolab Weather Monash (Godahewa et al., 2021) Climate H T/T T/T P
@@
Dataset Source Domain Frequency Chronos TimesFM Moirai Moirai-MoE Time-GPT ForecastPFN MOMENT Lag-Llama GPT4TS Timer ZSisionTS TTM LLMTime Time-LMM UniTime Time-MoE
Pedestrian Counts Monash (Godahewa et al., 2021) Transport H T/T ZS T/T T/T P ZS P ZS ZS P
Rideshare Monash (Godahewa et al., 2021) Transport H T/T T/T T/T P ZS P
Saugeen Monash (Godahewa et al., 2021) Nature D ZS T/T T/T P P ZS P ZS P
Saugeen Monash (Godahewa et al., 2021) Nature W-THU
Saugeen Monash (Godahewa et al., 2021) Nature M
Solar Power (Australia) Monash (Godahewa et al., 2021) Energy 4S T/T T/T P P
Solar Monash (Godahewa et al., 2021) Energy W ZS P P ZS
Solar Monash (Godahewa et al., 2021) Energy 10M P P P
Sunspot Monash (Godahewa et al., 2021) Nature D ZS T/T T/T P P P ZS P ZS P
Temperature Rain Monash (Godahewa et al., 2021) Nature D T/T T/T T/T P P ZS P
Tourism Monthly Monash (Godahewa et al., 2021) Econ/Fin M ZS ZS P T/T ZS ZS P
Tourism Quarterly Monash (Godahewa et al., 2021) Econ/Fin Q ZS ZS P T/T ZS ZS P
Tourism Yearly Monash (Godahewa et al., 2021) Econ/Fin Y ZS ZS P P P T/T ZS P
Traffic Hourly Monash (Godahewa et al., 2021) Transport H ZS P T/T T/T T/T P T/T ZS ZS ZS ZS T/T P
Traffic Weekly Monash (Godahewa et al., 2021) Transport W ZS T/T T/T ZS P ZS ZS P
US Births Monash (Godahewa et al., 2021) Healthcare D ZS T/T T/T P P ZS P ZS P
US Births Monash (Godahewa et al., 2021) Healthcare W-TUE
US Births Monash (Godahewa et al., 2021) Healthcare M
Vehicle Trips Monash (Godahewa et al., 2021) Transport D T/T T/T P ZS P
Weather (Australia) Monash (Godahewa et al., 2021) Climate D ZS ZS T/T T/T P ZS ZS P T/T T/T P
Wind Farms Monash (Godahewa et al., 2021) Energy T T/T T/T P P
Wind Farms Monash (Godahewa et al., 2021) Energy H P P
Wind Farms Monash (Godahewa et al., 2021) Energy D P
Wind Farms Monash (Godahewa et al., 2021) Energy 5T P
Wind Power Monash (Godahewa et al., 2021) Energy 4S T/T T/T P P
Kaggle Web Traffic Daily Monash (Godahewa et al., 2021) Web D P
Restaurant Recruit Other Sales D
Subseasonal Other Climate D P P P
Subseasonal Precipitation Other Climate D P P P
USHCN Other Weather D P
Mexico City Bikes Other Transport H P P
Dominick Other Sales D ZS P P
ERCOT Load Other Energy H ZS P
@@
TDBrain Other Healthcare 0.002S P
@@
BDG-2 Bull ProEnFo (Wang et al., 2023b) Energy H P P P
BDG-2 Cockatoo ProEnFo (Wang et al., 2023b) Energy H P P P
BDG-2 Hog ProEnFo (Wang et al., 2023b) Energy H P P P
Covid19 Energy ProEnFo (Wang et al., 2023b) Energy H P P P
ELF ProEnFo (Wang et al., 2023b) Energy H P P
GEF12 ProEnFo (Wang et al., 2023b) Energy H P P P
GEF14 ProEnFo (Wang et al., 2023b) Energy H P P P
GEF17 ProEnFo (Wang et al., 2023b) Energy H P P
PDB ProEnFo (Wang et al., 2023b) Energy H P P P
Spanish ProEnFo (Wang et al., 2023b) Energy H P P P P
@@
Dataset Source Domain Frequency Chronos TimesFM Moirai Moirai-MoE Time-GPT ForecastPFN MOMENT Lag-Llama GPT4TS Timer ZSisionTS TTM LLMTime Time-LMM UniTime Time-MoE
@@
Electricity UCI ML Archive (Trindade, 2015) Energy 15T T/T P T/T P
Electricity UCI ML Archive (Trindade, 2015) Energy H T/T P ZS ZS T/T P T/T ZS ZS ZS ZS T/T T/T
Electricity UCI ML Archive (Trindade, 2015) Energy D ZS
Electricity UCI ML Archive (Trindade, 2015) Energy W-FRI T/T
Electricity UCI ML Archive (Trindade, 2015) Energy M T/T
Electricity UCI ML Archive (Trindade, 2015) Energy W P P
@@
SelfRegulationSCP1 UCR Archive (Ismail Fawaz et al., 2019) Healthcare - P P
@@
Healthcare5000 UCR Archive (Ismail Fawaz et al., 2019) Healthcare - T/T
@@
Worms UCR Archive (Ismail Fawaz et al., 2019) Nature - P
@@
AtrialFibrillation UCR Archive (Ismail Fawaz et al., 2019) Healthcare - P P
@@
Phoneme UCR Archive (Ismail Fawaz et al., 2019) Audio - P P
@@
SelfRegulationSCP2 UCR Archive (Ismail Fawaz et al., 2019) Healthcare - P P
@@
StarLightCurves UCR Archive (Ismail Fawaz et al., 2019) Nature - P P
@@
Dataset Source Domain Frequency Chronos TimesFM Moirai Moirai-MoE Time-GPT ForecastPFN MOMENT Lag-Llama GPT4TS Timer ZSisionTS TTM LLMTime Time-LMM UniTime Time-MoE
@@
EigenWorms UCR Archive (Ismail Fawaz et al., 2019) Nature - P P
@@
PigArtPressure UCR Archive (Ismail Fawaz et al., 2019) Nature - P P
@@
MotorImagery UCR Archive (Ismail Fawaz et al., 2019) Healthcare - P P
@@
Dataset Source Domain Frequency Chronos TimesFM Moirai Moirai-MoE Time-GPT ForecastPFN MOMENT Lag-Llama GPT4TS Timer ZSisionTS TTM LLMTime Time-LMM UniTime Time-MoE
@@
PigCVP UCR Archive (Ismail Fawaz et al., 2019) Healthcare - P P
@@
Weatherbench (Daily) Weather Bench Weather D P P
Weatherbench (Hourly) Weather Bench Weather H P P
Weatherbench (Weekly) Weather Bench Weather W P P
Wiki hourly Wikipedia Metrics Web H P
Wiki daily Wikipedia Metrics Web D P P P
Wiki weekly Wikipedia Metrics Web W P
Wiki monthly Wikipedia Metrics Web M P

### v2

@@ Appendix D TSFM Data Lineage
LargeST LargeST (Liu et al., 2023a) Transport 5T P P P P P P P P P P
Beijing Subway LibCity (Wang et al., 2023a) Transport 30T P P P P P P P P P
HZMetro LibCity (Wang et al., 2023a) Transport 15T P P P P P P P P P
Loop Seattle LibCity (Wang et al., 2023a) Transport 5T P P P ZS P T/T ZS
Loop Seattle LibCity (Wang et al., 2023a) Transport H ZS T/T ZS
Loop Seattle LibCity (Wang et al., 2023a) Transport D ZS T/T ZS
Los-Loop LibCity (Wang et al., 2023a) Transport 5T P P P P P P P T/T ZS P
M_DENSE LibCity (Wang et al., 2023a) Transport H ZS T/T ZS
M_DENSE LibCity (Wang et al., 2023a) Transport D ZS ZS T/T ZS
PEMS Bay LibCity (Wang et al., 2023a) Transport 5T P P P P P P P P P P P P
PEMS03 LibCity (Wang et al., 2023a) Transport 5T P P ZS P P P P P P P P P P
PEMS04 LibCity (Wang et al., 2023a) Transport 5T P P ZS P P P P P P P P P P
PEMS07 LibCity (Wang et al., 2023a) Transport 5T P P ZS P P P P P P P P P P
PEMS08 LibCity (Wang et al., 2023a) Transport 5T P P T/T ZS P P P P P P P P P P
Q-Traffic LibCity (Wang et al., 2023a) Transport 15T P P P P P P P P P P P P
SHMetro LibCity (Wang et al., 2023a) Transport 15T P P P P P P P P P P P
SZ-Taxi LibCity (Wang et al., 2023a) Transport 15T P P P ZS T/T ZS P
SZ-Taxi LibCity (Wang et al., 2023a) Transport H P ZS T/T ZS
Rotterdam LibCity (Wang et al., 2023a) Transport 2T P P P T/T ZS
Beijing Air Quality LOTSA_Others (Woo et al., 2024) Nature H P P P P P P P P P P P P P P
CDC Fluview ILINet LOTSA_Others (Woo et al., 2024) Healthcare W P P ZS T/T T/T ZS T/T T/T P P P P P P P P
CDC Fluview WHO NREVSS LOTSA_Others (Woo et al., 2024) Healthcare W P P ZS T/T T/T P P P P P P P P
China Air Quality LOTSA_Others (Woo et al., 2024) Nature H P P P P P P P P P P P
Favorita Sales LOTSA_Others (Woo et al., 2024) Sales D P P P P P P P P P P P
Favorita Transactions LOTSA_Others (Woo et al., 2024) Sales D P P P P P P P
GoDaddy LOTSA_Others (Woo et al., 2024) Econ/Fin M P P P P P P P P P
KDD Cup 2022 LOTSA_Others (Woo et al., 2024) Energy 10T P P P P P P P P P P
Project Tycho LOTSA_Others (Woo et al., 2024) Healthcare W P P P P P P P
Residential Load Power LOTSA_Others (Woo et al., 2024) Energy T P P P P P P P P P P
Residential PV Power LOTSA_Others (Woo et al., 2024) Energy T P P P P P P P P P P
@@
Solar LSTNet (Lai et al., 2017) Energy 10T ZS T/T T/T ZS T/T
Solar LSTNet (Lai et al., 2017) Energy 5T P possible P P P P
Solar LSTNet (Lai et al., 2017) Energy H P ZS ZS ZS ZS P P T/T ZS P
Solar LSTNet (Lai et al., 2017) Energy D ZS T/T ZS
Solar LSTNet (Lai et al., 2017) Energy W-FRI ZS T/T ZS
Exchange Rate LSTNet (Lai et al., 2017) Econ/Fin D ZS ZS T/T ZS ZS T/T P T/T ZS ZS
Hierarchical Sales (M5 Competition) Mancuso et al. (2020) Sales D P ZS T/T ZS
Hierarchical Sales (M5 Competition) Mancuso et al. (2020) Sales W-WED ZS T/T ZS
Australian Electricity Demand Monash (Godahewa et al., 2021) Energy 30T ZS ZS T/T T/T P P P ZS P ZS P P P T/T P P ZS ZS P
Bitcoin Monash (Godahewa et al., 2021) Econ/Fin D ZS T/T T/T P P ZS P ZS P P P T/T P P P
Car Parts Monash (Godahewa et al., 2021) Sales M T/T T/T P ZS P ZS T/T T/T ZS ZS ZS
CIF 2016 Monash (Godahewa et al., 2021) Econ/Fin M ZS ZS T/T T/T P ZS ZS P P T/T P P ZS ZS
COVID Deaths Monash (Godahewa et al., 2021) Healthcare D ZS ZS T/T T/T P ZS ZS P ZS T/T T/T ZS ZS ZS
Covid Mobility Monash (Godahewa et al., 2021) Transport D T/T T/T P P T/T P P
Elecdemand Monash (Godahewa et al., 2021) Energy 30T T/T T/T P P T/T P P
Extended Web Traffic Monash (Godahewa et al., 2021) Web D T/T T/T P P P P T/T P P
FRED MD Monash (Godahewa et al., 2021) Econ/Fin M ZS ZS T/T T/T P ZS ZS P P P T/T P P ZS ZS P
Hospital Monash (Godahewa et al., 2021) Healthcare M ZS ZS T/T T/T P ZS ZS P ZS T/T T/T ZS ZS ZS
Kaggle Web Traffic Weekly Monash (Godahewa et al., 2021) Web W P P P P P P P P P P P
KDD Cup 2018 Monash (Godahewa et al., 2021) Nature H T/T T/T T/T P P P ZS P P ZS T/T T/T T/T ZS P
KDD Cup 2018 Monash (Godahewa et al., 2021) Nature D ZS P T/T ZS P
London Smart Meters Monash (Godahewa et al., 2021) Energy 30T T/T P P P P P P P P P P T/T P P P P P
M1 Monthly Monash (Godahewa et al., 2021) Econ/Fin M ZS T/T T/T P ZS P P T/T P P ZS ZS
M1 Quarterly Monash (Godahewa et al., 2021) Econ/Fin Q ZS P P P P P P P P ZS ZS
M1 Yearly Monash (Godahewa et al., 2021) Econ/Fin Y ZS P P P P P P P P ZS ZS
M3 Monthly Monash (Godahewa et al., 2021) Econ/Fin M ZS T/T T/T ZS T/T ZS P P T/T P P ZS ZS
M3 Other Monash (Godahewa et al., 2021) Econ/Fin Q T/T T/T T/T ZS P T/T P P
M3 Quarterly Monash (Godahewa et al., 2021) Econ/Fin Q ZS P P ZS T/T P P P P P ZS ZS
M3 Yearly Monash (Godahewa et al., 2021) Econ/Fin Y ZS P P ZS T/T P P P P P ZS ZS
M4 Daily Monash (Godahewa et al., 2021) Econ/Fin D T/T P T/T T/T P T/T ZS P ZS T/T T/T T/T ZS P P
M4 Hourly Monash (Godahewa et al., 2021) Econ/Fin H T/T P T/T T/T P T/T ZS P ZS T/T T/T T/T ZS P P
M4 Monthly Monash (Godahewa et al., 2021) Econ/Fin M T/T P T/T T/T ZS T/T ZS P ZS T/T T/T T/T ZS P P
M4 Quarterly Monash (Godahewa et al., 2021) Econ/Fin Q-DEC ZS P P P ZS T/T P ZS P T/T ZS ZS P
M4 Weekly Monash (Godahewa et al., 2021) Econ/Fin W-SUN T/T P T/T T/T T/T ZS P ZS T/T T/T T/T ZS P
M4 Yearly Monash (Godahewa et al., 2021) Econ/Fin A-DEC ZS P P P ZS T/T P ZS P T/T ZS ZS P
NN5 Daily Monash (Godahewa et al., 2021) Econ/Fin D ZS ZS T/T T/T P ZS P ZS P P P T/T P P ZS ZS P
NN5 Weekly Monash (Godahewa et al., 2021) Econ/Fin W ZS ZS T/T T/T P ZS ZS P T/T P P ZS ZS
Oikolab Weather Monash (Godahewa et al., 2021) Climate H T/T T/T P P T/T P P
Pedestrian Counts Monash (Godahewa et al., 2021) Transport H T/T ZS T/T T/T P ZS P ZS ZS P P P T/T T/T P P P P
Rideshare Monash (Godahewa et al., 2021) Transport H T/T T/T T/T P ZS P P T/T T/T P P P
@@
Dataset Source Domain Frequency Chronos TimesFM Moirai Moirai-MoE Time-GPT ForecastPFN MOMENT Lag-Llama GPT4TS Timer VisionTS TTM LLMTime Time-LMM UniTime Time-MoE Chronos-2 Flowstate(under review) LightGTS Sundial TabPFN-TS Time-MoE2 TinyTimeMixer TiRex Toto YingLong (under review) Gift-Eval TOTO2 Flowstate(T) TimesFM-ICL ROSE
Saugeen Monash (Godahewa et al., 2021) Nature D ZS T/T T/T P P ZS P ZS P ZS P T/T T/T ZS P
Saugeen Monash (Godahewa et al., 2021) Nature W-THU ZS T/T ZS
Saugeen Monash (Godahewa et al., 2021) Nature M ZS T/T ZS
Solar Power (Australia) Monash (Godahewa et al., 2021) Energy 4S T/T T/T P P P P T/T P P P P P
Solar Monash (Godahewa et al., 2021) Energy W ZS P P ZS ZS
Solar Monash (Godahewa et al., 2021) Energy 10M P P P ZS
Sunspot Monash (Godahewa et al., 2021) Nature D ZS T/T T/T P P P ZS P ZS P P P T/T P P P
Temperature Rain Monash (Godahewa et al., 2021) Nature D T/T T/T T/T P P ZS P ZS P T/T T/T T/T ZS P P
Tourism Monthly Monash (Godahewa et al., 2021) Econ/Fin M ZS ZS P T/T ZS ZS P P P P ZS ZS
Tourism Quarterly Monash (Godahewa et al., 2021) Econ/Fin Q ZS ZS P T/T ZS ZS P P P P ZS ZS
Tourism Yearly Monash (Godahewa et al., 2021) Econ/Fin Y ZS ZS P P P T/T ZS P P P P P ZS ZS
Traffic Hourly Monash (Godahewa et al., 2021) Transport H ZS P T/T T/T T/T P T/T ZS ZS ZS ZS T/T P P T/T T/T P P ZS P T/T
Traffic Weekly Monash (Godahewa et al., 2021) Transport W ZS T/T T/T ZS P ZS ZS P P T/T P P
US Births Monash (Godahewa et al., 2021) Healthcare D ZS T/T T/T P P ZS P ZS P ZS P T/T T/T P
US Births Monash (Godahewa et al., 2021) Healthcare W-TUE ZS T/T ZS
US Births Monash (Godahewa et al., 2021) Healthcare M ZS T/T ZS
Vehicle Trips Monash (Godahewa et al., 2021) Transport D T/T T/T P ZS P P T/T P P
Weather (Australia) Monash (Godahewa et al., 2021) Climate D ZS ZS T/T T/T P ZS ZS P T/T T/T P P P T/T P P ZS P
Wind Farms Monash (Godahewa et al., 2021) Energy T T/T T/T P P P P T/T P P P
Wind Farms Monash (Godahewa et al., 2021) Energy H P P P P P
Wind Farms Monash (Godahewa et al., 2021) Energy D P possible P P P P
Wind Farms Monash (Godahewa et al., 2021) Energy 5T P possible P
Wind Power Monash (Godahewa et al., 2021) Energy 4S T/T T/T P P P P T/T P P P P P P
Kaggle Web Traffic Daily Monash (Godahewa et al., 2021) Web D P P P P
Restaurant Recruit Other Sales D ZS T/T ZS
Subseasonal Other Climate D P P P P P P P P P P
Subseasonal Precipitation Other Climate D P P P P P P P P P
USHCN Other Weather D P possible P P P P P
Mexico City Bikes Other Transport H P P possible P P P P P
Dominick Other Sales D ZS P P P ZS ZS
ERCOT Load Other Energy H ZS P P ZS ZS
@@
TDBrain Other Healthcare 0.002S P P P
@@
BDG-2 Bull ProEnFo (Wang et al., 2023b) Energy H P P P P P P P
BDG-2 Cockatoo ProEnFo (Wang et al., 2023b) Energy H P P P P P P P
BDG-2 Hog ProEnFo (Wang et al., 2023b) Energy H P P P P P P P
Covid19 Energy ProEnFo (Wang et al., 2023b) Energy H P P P P P P P
ELF ProEnFo (Wang et al., 2023b) Energy H P P P P P P
GEF12 ProEnFo (Wang et al., 2023b) Energy H P P P P P P P
GEF14 ProEnFo (Wang et al., 2023b) Energy H P P P P P P P
GEF17 ProEnFo (Wang et al., 2023b) Energy H P P P P P P
PDB ProEnFo (Wang et al., 2023b) Energy H P P P P P P P
Spanish ProEnFo (Wang et al., 2023b) Energy H P P P P P P P P P P
@@
Electricity UCI ML Archive (Trindade, 2015) Energy 15T T/T P T/T P ZS T/T T/T T/T ZS P
Electricity UCI ML Archive (Trindade, 2015) Energy H T/T P ZS ZS T/T P T/T ZS ZS ZS ZS T/T T/T ZS T/T ZS T/T T/T ZS P P T/T
Electricity UCI ML Archive (Trindade, 2015) Energy D ZS ZS T/T ZS
Electricity UCI ML Archive (Trindade, 2015) Energy W-FRI T/T ZS T/T T/T ZS
Electricity UCI ML Archive (Trindade, 2015) Energy M T/T T/T
Electricity UCI ML Archive (Trindade, 2015) Energy W P P P
@@
SelfRegulationSCP1 UCR Archive (Ismail Fawaz et al., 2019) Healthcare - P P P P
@@
Healthcare5000 (orignally: ECG5000) UCR Archive (Ismail Fawaz et al., 2019) Healthcare - T/T P
@@
Worms UCR Archive (Ismail Fawaz et al., 2019) Nature - P P P
@@
AtrialFibrillation UCR Archive (Ismail Fawaz et al., 2019) Healthcare - P P P P
@@
Phoneme UCR Archive (Ismail Fawaz et al., 2019) Audio - P P P P
@@
SelfRegulationSCP2 UCR Archive (Ismail Fawaz et al., 2019) Healthcare - P P P P
@@
StarLightCurves UCR Archive (Ismail Fawaz et al., 2019) Nature - P P P P
@@
Dataset Source Domain Frequency Chronos TimesFM Moirai Moirai-MoE Time-GPT ForecastPFN MOMENT Lag-Llama GPT4TS Timer VisionTS TTM LLMTime Time-LMM UniTime Time-MoE Chronos-2 Flowstate(under review) LightGTS Sundial TabPFN-TS Time-MoE2 TinyTimeMixer TiRex Toto YingLong (under review) Gift-Eval TOTO2 Flowstate(T) TimesFM-ICL ROSE
@@
EigenWorms UCR Archive (Ismail Fawaz et al., 2019) Nature - P P P P
@@
PigArtPressure UCR Archive (Ismail Fawaz et al., 2019) Nature - P P P P
@@
MotorImagery UCR Archive (Ismail Fawaz et al., 2019) Healthcare - P P P P
@@
PigCVP UCR Archive (Ismail Fawaz et al., 2019) Healthcare - P P P P
@@
Weatherbench (Daily) Weather Bench Weather D P P possible P P P P P
Weatherbench (Hourly) Weather Bench Weather H P P possible P P P P P
Weatherbench (Weekly) Weather Bench Weather W P P possible P P P P P
Wiki hourly Wikipedia Metrics Web H P P
Wiki daily Wikipedia Metrics Web D P P P possible P P P P P
Wiki weekly Wikipedia Metrics Web W P P
Wiki monthly Wikipedia Metrics Web M P P
@@
ERA5 Munoz-Sabater et al., 2021 Weather 3h P P
ERA5 Munoz-Sabater et al., 2021 Weather 12h P
ERA5 Munoz-Sabater et al., 2021 Weather D P
ERA5 Munoz-Sabater et al., 2021 Weather W P
ERA5 Munoz-Sabater et al., 2021 Weather M P
ERA5 Munoz-Sabater et al., 2021 Weather Q P
Finance (not published) Liu et al., 2025 (Sundial) Finance ? P
IoT (not published) Liu et al., 2025 (Sundial) IoT ? P
BOOM Cohen et al. (2025) (TOTO) Observational 10S ZS
BOOM Cohen et al. (2025) (TOTO) Observational T ZS
BOOM Cohen et al. (2025) (TOTO) Observational 5T ZS
BOOM Cohen et al. (2025) (TOTO) Observational 10T ZS
BOOM Cohen et al. (2025) (TOTO) Observational 30T ZS
BOOM Cohen et al. (2025) (TOTO) Observational H ZS
BOOM Cohen et al. (2025) (TOTO) Observational D